What is the molality of a solution formed from 29.1 g of C7H8 (toluene) dissolved in a 832 g of benzene (C6H6)?

1
Expert's answer
2022-03-16T13:52:03-0400

Mass of toluene = 29.1 g , mass of solvent = 832 g


Molar mass of "C_7H_8" = 84+ 8= 92 "\frac{g}{mol}"


Molality = "\frac{29.1 \times 1000}{92 \times 832}= 0.38" m